# Small-Signal Stability Oriented Real-Time Operation of Power Systems with a High Penetration of Inverter-Based Resources

### v1

@@ I-A Context and Motivation
The increasing penetration of renewable energy sources (RES), combined with the progressive phase-out of fossil-fuel-based synchronous generation, is transforming the electrical grid into a power-electronics-dominated system. This transition is profoundly changing the grid’s dynamic behavior: traditional synchronous generators (SGs) inherently provide inertia and damping to electromechanical oscillations, whereas inverter-based resources (IBRs) exhibit fast dynamics governed by their internal control loops. As a result, modern power systems exhibit reduced inertia and are dominated by faster, control-driven phenomena occurring on the timescale of converter dynamics [StabilityTop, milano2018foundations]. Moreover, for both environmental and economic reasons, system operation is incentivized to maximize the utilization of renewable generation, particularly from non-dispatchable sources such as photovoltaic (PV) and wind power. This trend further pushes power systems to operate closer to their stability limits, where small disturbances or unforeseen fluctuations in generation and demand can lead to dynamic instabilities. Recent European network blackout events, such as those that occurred in the Iberian Peninsula and North Macedonia, serve as illustrative examples. Although investigations into the root causes of these events are ongoing, early analyses suggest that they were not caused by physical damage to infrastructure or extreme weather conditions, but rather occurred under normal operating conditions — one during periods of high renewable generation [entsoe2025_iberian_factual], and another during low nighttime weekend demand [entsoe2025_northMacedonia_incident]. These events have highlighted that transmission system operators (TSOs) currently lack adequate real-time tools to monitor and manage stability margins under rapidly changing operating conditions.
@@
To address these challenges, numerous studies have proposed incorporating stability constraints, both static and dynamic, into the optimal power flow (OPF) problem. Following the conventional classification of power system stability [StabilityTop], different stability-constrained OPF formulations can be defined. Transient, voltage, and frequency stability-constrained OPF addresses the effects of large disturbances, providing the optimal pre-contingency generators dispatch able to withstand contingencies as faults and outages [book_scopt]. When considering small perturbations, small-signal stability-constrained OPF (SSSC-OPF) can be employed. However, if the small-signal model is properly formulated to capture electromagnetic transient (EMT) phenomena, SSSC-OPF is not limited to prevent instabilities due to small-disturbances, but can also detect and prevent from converter-driven and resonance-induced instabilities. Consequently, the SSSC-OPF formulation ensures that the system does not exhibit undamped oscillations during normal operation or under small disturbances.
@@
Concerning the real-time implementation of optimal operation algorithms, Online Feedback Optimization (OFO) has emerged as a promising approach. The fundamental concept of OFO is the implementation of optimization algorithms as feedback controllers, which are coupled to physical plants via a closed-loop scheme [Florian2020]. Through this interconnection, the controller drives the operating point towards the solution of the optimization problem using real-time measurements instead of computing the equations that describe the system’s behavior. As a result, the system exhibits robustness to disturbances and uncertainties. Furthermore, by circumventing the numerical evaluation of the nonlinear plant model, the computational burden is significantly reduced while preserving the industrial property [olives2025uac].
@@ I-B1 Small-signal Stability-Constrained Optimal Power Flow
Numerous formulations have been proposed in the literature to improve the tractability of the SSSC-OPF problem by circumventing the explicit inclusion of the full set of differential algebraic equations (DAE) describing system dynamics [stabilityconstrained, Milano2011, thams2017data]. Some approaches rely on the linearized state-space representation of these DAE and the corresponding eigenvalue analysis, as commonly employed in small-signal stability assessment [kundur]. However, the relationship between the eigenvalue-based indicators (e.g., damping ratios and real parts of eigenvalues) and the system states and parameters is implicit [SSSCOPF_SVM_2022]. To address this issue, several studies have proposed sensitivity-based formulations, where the stability constraint is expressed in terms of the linear sensitivities of the system state matrix eigenvalues with respect to small variations in generators’ active power [Milano2011, 7539295]. Despite limiting the analysis to a number of critical eigenvalues, the recalculation of the sensitivities at each iteration remains a necessity, making these approaches unsuitable for real-time operation.
@@
The employment of machine learning (ML)-based surrogate models can facilitate the formulation of stability constraints, thereby alleviating the computational burden [hasan2020survey]. This approach, generally named Constraint Learning (CL) [maragno2023mixed, fajemisin2024optimization], is particularly effective when the analytical expression of the intended constraint is unknown or complex [fajemisin2024optimization]. Moreover, ML techniques shift the most computationally demanding tasks to the offline phase, during which the model is trained and the necessary training data are generated and collected. Once trained, the model can be used online to provide a fast and efficient evaluation of the stability constraint. As reviewed in [maragno2023mixed], commonly used ML models for CL include both linear approaches, such as linear regression and Support Vector Machines (SVMs) with linear kernels, and nonlinear models, including Decision Trees (DTs), ensemble methods, and Neural Networks (NNs). While these nonlinear models can capture the behavior of small-signal stability indices, their integration into OPF formulations typically requires linearization techniques involving binary variables [maragno2023mixed], as demonstrated in [9141308, thams2017data]. This results in mixed-integer nonlinear programming (MINLP) problems, whose solution for large-scale power systems remains computationally challenging and often intractable [capitanescu2011state]. In [liu2021explicit], an SVM is employed to formulate a linear small-signal stability constraint with good computational efficiency. However, the method relies on locally trained models built around each unstable operating point, making it suitable for corrective re-dispatch rather than real-time preventive control.
@@
In a previous work by the authors [rossiSSSCOPF2024], the performance of linear models was compared against Multivariate Adaptive Regression Splines (MARS) [friedman1991multivariate], a nonlinear regression technique identified as suitable for inclusion in OPF formulations. MARS provides a continuous and differentiable analytical expression (class $\mathcal{C}^{1}$[friedman1991multivariate], making it compatible with OPF problems solved via sequential quadratic programming (SQP), in contrast to the previously mentioned nonlinear models. Moreover, MARS demonstrated superior robustness, with respect to the linear models, in estimating small-signal stability indicators across OPF formulations with different objective functions.
@@ I-B2 Online Feedback Optimization in Power Systems Operation
Up to now, OFO has been mainly investigated in the context of distribution system operation, where its advantages are particularly significant due to the lack of accurate grid models and the uncertainty associated with RES. Applications in the literature include voltage regulation [ORTMANN2020106782, olives2022ac] and optimal reactive power dispatch [6930779]. More recently, its applicability to transmission systems has also been demonstrated, where its main advantages lie in the smooth, continuous control actions it enables against disturbances. Although TSOs typically possess system models, their accuracy may be limited, and relying directly on real-time measurements can offer a more robust and adaptive alternative [ofo_tso_dso_2025]. Furthermore, incorporating small-signal stability constraints into OFO can also be relevant for distribution systems operating in islanded or weakly connected modes, provided that suitable small-signal models (obtained through dynamic system identification or black-box modeling) are available.
@@ II-A Small-Signal Stability Constraint Formulation
To formulate a constraint on system dynamics for inclusion in the optimal operation problem, a regression-based approach is employed. The use of regression provides a compact surrogate model of an otherwise complex problem. The regression-based surrogate is trained to provide the same outcome as the conventional eigenvalue-based analysis of the linearized state-space of the system, while offering a significantly simpler formulation that reduces the dimensionality of the problem. Moreover, regression techniques provide flexibility in selecting the mathematical representation most suitable depending on how the optimization problem is formulated. In this work, MARS is employed for its ability and robustness in capturing the nonlinear relation between the OPF control variables and the stability indicator [rossiSSSCOPF2024], and for providing an analytic expression of class $\mathcal{C}^{1}$, suitable for formulating a smooth constraint in an SQP-based optimization.
To develop the regression model, a representative training dataset is required. This dataset must capture the mapping between system operating conditions and small-signal stability across the entire operable space, while providing finer granularity near the stability boundary — i.e., the region where both stable and unstable operating points occur. A dataset with such features needs to be generated synthetically [venzke2021efficient], by performing stability analyses for a large number of operating points sampled within the operable space. Formally, the training dataset can be expressed as $\mathcal{D}=(\mathbf{X}\mid\mathbf{z})$. The matrix $\mathbf{X}$ contains the system state variables for each sampled operating point, including bus voltage magnitudes and phases, as well as generators and loads active and reactive power injections. The vector $\mathbf{z}$ stores the corresponding stability assessment outcomes. Since regression requires a continuous numerical target, $\mathbf{z}$ is expressed in terms of a stability index. For small-signal stability, this can be the Damping Index (DI) defined in [collados2022stability], given as
@@
where $\xi_{i}$ denotes the damping ratio of the $i$-th complex eigenvalue $\lambda_{i}$, and $|\mathbf{\lambda}|$ is the number of critical eigenvalues, as defined in [rossiML2022]. Given this formulation of the DI, small-signal stability is satisfied if $\text{DI}<1$. Moreover, a stricter requirement can be imposed by introducing a threshold $\theta<1$, such that $\text{DI}<\theta$, thereby enforcing a prescribed damping margin.
@@
where $t$ is a point within the range of $\mathbf{x}_{i}$ in the training dataset, referred to as a knot. Although each basis function depends only on the single variable $\mathbf{x}_{i}$, it is formally regarded as a function defined over the entire input space [hastie2009elements]. During training, the MARS algorithm automatically determines the placement of knots and selects the hinge functions that contribute most to improving regression accuracy while ensuring generalization. The resulting regression model takes the form
@@ II-B Online Feedback Optimization
The OFO technique is based on the integration of the dynamics of a numerical optimization solver and the system dynamics to be controlled within a closed-loop structure. In this configuration, the solver is employed to update the control inputs for the system in response to the measurements obtained from the system. Note that this approach does not require an accurate plant model since is based on real-time measurements. Instead, it only requires the system’s input-output sensitivities, which can be obtained on site [olives2022ac]. Therefore, the performance of the controller remains unaffected by model inaccuracies and is robust to disturbances and uncertainty, as these effects are implicit in the measurements.
The objective of OFO-based techniques is to steer the power system towards a steady-state optimal operating point. Therefore, under certain assumptions concerning the separation of time scales and the rapid mitigation of transients [OFO_stability2018], the power system may be solely defined by its steady-state behavior. In this context, consider a set of nonlinear algebraic equations $\mathbf{f}(\cdot)$, as follows:
@@
The nonconvexity of the optimization problem (5a)-(5d) is addressed through the implementation of a projected gradient flow algorithm as presented in [haberle2020non]. Starting from the control action $\mathbf{u}$ in the current iteration, the control algorithm proceeds to update the value ($\mathbf{u}^{+}$) using a projected gradient step method with a fixed step size [^1] A fixed step size algorithm must be used as an adverse consequence of not evaluating the set of equations $\mathbf{f}(\cdot)$ numerically. as follows:
@@
$\hat{\boldsymbol{\sigma}}(\mathbf{u,y}):=\arg\min_{\boldsymbol{\delta}\in\mathbb{R}^{p}}$ $\;\|\boldsymbol{\delta}+\gamma\mathbf{G}^{-1}\mathbf{H}^{\top}\nabla f(\mathbf{y,u})^{\top}\|_{\mathbf{G}}^{2}$ (8a)
@@ II-C Small-signal Stability Constrained Online Feedback Optimization
$\underset{\mathbf{u}}{\operatorname{minimize}}\quad$ $\gamma\cdot f(\mathbf{y},\mathbf{u})$
@@ III-A Test System
The simulation test bench is based on a modified version of the IEEE 9-bus system [anderson2003power]. It represents a three-phase high-voltage transmission network consisting of nine buses, six transmission lines, three transformers, three load substations, and three generators. In contrast to the original system, which includes three SGs, two SGs are replaced with IBRs to emulate a scenario with high penetration of power electronics devices. Furthermore, to account for the impact of control strategies on system stability, one IBR is operated under GFM control while the other is operated under GFL control.
Figure 1 shows the one-line diagram of the system under analysis. The electrical lines of the system are considered symmetrical and are modeled using a $\pi$-model. The character of the lines is predominantly inductive, and all parameters are obtained from [anderson2003power]. Power substations are modeled as constant impedance balanced loads (R-L shunts load model). Table I summarizes the main characteristics of the substations demand, including the load participation factors (i.e., the percentage of the total system demand allocated to each load), the load power factors, and the overall range of system demand. Concerning the generators, it is assumed that both SG and IBRs are equipped with a primary power source that allows them to operate effectively at any feasible operating point. In addition, all generators have a communication system with a centralized control layer that calculates their setpoints. At the local level, the SG is incorporated with a governor and an AVR to regulate its steady-state outputs to the values dispatched by the proposed control layer. Therefore, the setpoints for SGs are defined as the injected active power and the generated voltage amplitude, measured at the bus to which they are connected. The IBR connected to bus 1 is configured to operate in a GFM mode, and therefore indicated as $\text{GFM}_{1}$. The device is equipped with an internal current control loop and an external voltage control loop. In addition, it contains an active power droop control stage, employed for synchronization, and reactive power droop. The IBR connected to bus 3 is configured to operate in a GFL mode, and therefore indicated as $\text{GFL}_{3}$. The GFL control scheme comprises a Phase-Locked Loop (PLL), an inner current control loop, an outer control loop for active power fed by a frequency droop control, and an outer reactive power control fed by a voltage droop. Table II summarizes the main generator parameters, including their rated power, rated voltage, and the assumed operating active power ranges.

### v2

@@ I-A Context and Motivation
The increasing penetration of renewable energy sources (RES), combined with the progressive phase-out of fossil-fuel-based synchronous generation, is transforming the electrical grid into a power-electronics-dominated system. This transition is profoundly changing the grid’s dynamic behavior: traditional synchronous generators (SGs) inherently provide inertia and damping to electromechanical oscillations, whereas inverter-based resources (IBRs) exhibit fast dynamics governed by their internal control loops. As a result, modern power systems exhibit reduced inertia and are dominated by faster, control-driven phenomena occurring on the timescale of converter dynamics [16, 25]. Moreover, for both environmental and economic reasons, system operation is incentivized to maximize the utilization of renewable generation, particularly from non-dispatchable sources such as photovoltaic (PV) and wind power. This trend further pushes power systems to operate closer to their stability limits, where small disturbances or unforeseen fluctuations in generation and demand can lead to dynamic instabilities. Recent European network blackout events, such as those that occurred in the Iberian Peninsula and North Macedonia, serve as illustrative examples. Although investigations into the root causes of these events are ongoing, early analyses suggest that they were not caused by physical damage to infrastructure or extreme weather conditions, but rather occurred under normal operating conditions — one during periods of high renewable generation [8], and another during low nighttime weekend demand [9]. These events have highlighted that transmission system operators (TSOs) currently lack adequate real-time tools to monitor and manage stability margins under rapidly changing operating conditions.
@@
To address these challenges, numerous studies have proposed incorporating stability constraints, both static and dynamic, into the optimal power flow (OPF) problem. Following the conventional classification of power system stability [16], different stability-constrained OPF formulations can be defined. Transient, voltage, and frequency stability-constrained OPF addresses the effects of large disturbances, providing the optimal pre-contingency generators dispatch able to withstand contingencies as faults and outages [35]. When considering small perturbations, small-signal stability-constrained OPF (SSSC-OPF) can be employed. However, if the small-signal model is properly formulated to capture electromagnetic transient (EMT) phenomena, SSSC-OPF is not limited to prevent instabilities due to small-disturbances, but can also detect and prevent from converter-driven and resonance-induced instabilities. Consequently, the SSSC-OPF formulation ensures that the system does not exhibit undamped oscillations during normal operation or under small disturbances.
@@
Concerning the real-time implementation of optimal operation algorithms, Online Feedback Optimization (OFO) has emerged as a promising approach. The fundamental concept of OFO is the implementation of optimization algorithms as feedback controllers, which are coupled to physical plants via a closed-loop scheme [7]. Through this interconnection, the controller drives the operating point towards the solution of the optimization problem using real-time measurements instead of computing the equations that describe the system’s behavior. As a result, the system exhibits robustness to disturbances and uncertainties. Furthermore, by circumventing the numerical evaluation of the nonlinear plant model, the computational burden is significantly reduced while preserving the industrial property [27].
@@ I-B1 Small-signal Stability-Constrained Optimal Power Flow
Numerous formulations have been proposed in the literature to improve the tractability of the SSSC-OPF problem by circumventing the explicit inclusion of the full set of differential algebraic equations (DAE) describing system dynamics [12, 36, 32]. Some approaches rely on the linearized state-space representation of these DAE and the corresponding eigenvalue analysis, as commonly employed in small-signal stability assessment [18]. However, the relationship between the eigenvalue-based indicators (e.g., damping ratios and real parts of eigenvalues) and the system states and parameters is implicit [21]. To address this issue, several studies have proposed sensitivity-based formulations, where the stability constraint is expressed in terms of the linear sensitivities of the system state matrix eigenvalues with respect to small variations in generators’ active power [36, 19]. Despite limiting the analysis to a number of critical eigenvalues, the recalculation of the sensitivities at each iteration remains a necessity, making these approaches unsuitable for real-time operation.
@@
The employment of machine learning (ML)-based surrogate models can facilitate the formulation of stability constraints, thereby alleviating the computational burden [14]. This approach, generally named Constraint Learning (CL) [22, 10], is particularly effective when the analytical expression of the intended constraint is unknown or complex [10]. Moreover, ML techniques shift the most computationally demanding tasks to the offline phase, during which the model is trained and the necessary training data are generated and collected. Once trained, the model can be used online to provide a fast and efficient evaluation of the stability constraint. As reviewed in [22], commonly used ML models for CL include both linear approaches, such as linear regression and Support Vector Machines (SVMs) with linear kernels, and nonlinear models, including Decision Trees (DTs), ensemble methods, and Neural Networks (NNs). While these nonlinear models can capture the behavior of small-signal stability indices, their integration into OPF formulations typically requires linearization techniques involving binary variables [22], as demonstrated in [33, 32]. This results in mixed-integer nonlinear programming (MINLP) problems, whose solution for large-scale power systems remains computationally challenging and often intractable [4]. In [20], an SVM is employed to formulate a linear small-signal stability constraint with good computational efficiency. However, the method relies on locally trained models built around each unstable operating point, making it suitable for corrective re-dispatch rather than real-time preventive control.
@@
In a previous work by the authors [31], the performance of linear models was compared against Multivariate Adaptive Regression Splines (MARS) [11], a nonlinear regression technique identified as suitable for inclusion in OPF formulations. MARS provides a continuous and differentiable analytical expression (class $\mathcal{C}^{1}$[11], making it compatible with OPF problems solved via sequential quadratic programming (SQP), in contrast to the previously mentioned nonlinear models. Moreover, MARS demonstrated superior robustness, with respect to the linear models, in estimating small-signal stability indicators across OPF formulations with different objective functions.
@@ I-B2 Online Feedback Optimization in Power Systems Operation
Up to now, OFO has been mainly investigated in the context of distribution system operation, where its advantages are particularly significant due to the lack of accurate grid models and the uncertainty associated with RES. Applications in the literature include voltage regulation [28, 26] and optimal reactive power dispatch [3]. More recently, its applicability to transmission systems has also been demonstrated, where its main advantages lie in the smooth, continuous control actions it enables against disturbances. In particular, feedback-based online optimization has been shown to provide stable tracking of optimal operating points in transmission networks, effectively combining frequency regulation and economic dispatch in a single real-time control loop [6, 23]. Large-scale studies on real subtransmission grids further show that OFO can enforce voltage and line-flow limits, handle discrete devices such as tap changers, and reduce renewable curtailment even under model uncertainty [29]. Although TSOs typically possess system models, their accuracy may be limited, and relying directly on real-time measurements can offer a more robust and adaptive alternative [17]. Furthermore, incorporating small-signal stability constraints into OFO can also be relevant for distribution systems operating in islanded or weakly connected modes, provided that suitable small-signal models (obtained through dynamic system identification or black-box modeling) are available.
@@ II-A Small-Signal Stability Constraint Formulation
To formulate a constraint on system dynamics for inclusion in the optimal operation problem, a regression-based approach is employed. The use of regression provides a compact surrogate model of an otherwise complex problem. The regression-based surrogate is trained to provide the same outcome as the conventional eigenvalue-based analysis of the linearized state-space of the system, while offering a significantly simpler formulation that reduces the dimensionality of the problem. Moreover, regression techniques provide flexibility in selecting the mathematical representation most suitable depending on how the optimization problem is formulated. In this work, MARS is employed for its ability and robustness in capturing the nonlinear relation between the OPF control variables and the stability indicator [31], and for providing an analytic expression of class $\mathcal{C}^{1}$, suitable for formulating a smooth constraint in an SQP-based optimization.
To develop the regression model, a representative training dataset is required. This dataset must capture the mapping between system operating conditions and small-signal stability across the entire operable space, while providing finer granularity near the stability boundary — i.e., the region where both stable and unstable operating points occur. A dataset with such features needs to be generated synthetically [34], by performing stability analyses for a large number of operating points sampled within the operable space. Formally, the training dataset can be expressed as $\mathcal{D}=(\mathbf{X}\mid\mathbf{z})$. The matrix $\mathbf{X}$ contains the system state variables for each sampled operating point, including bus voltage magnitudes and phases, as well as generators and loads active and reactive power injections. The vector $\mathbf{z}$ stores the corresponding stability assessment outcomes. Since regression requires a continuous numerical target, $\mathbf{z}$ is expressed in terms of a stability index. For small-signal stability, this can be the Damping Index (DI) defined in [5], given as
@@
where $\xi_{i}$ denotes the damping ratio of the $i$-th complex eigenvalue $\lambda_{i}$, and $|\mathbf{\lambda}|$ is the number of critical eigenvalues, as defined in [30]. Given this formulation of the DI, small-signal stability is satisfied if $\text{DI}<1$. Moreover, a stricter requirement can be imposed by introducing a threshold $\theta<1$, such that $\text{DI}<\theta$, thereby enforcing a prescribed damping margin.
@@
where $t$ is a point within the range of $\mathbf{x}_{i}$ in the training dataset, referred to as a knot. Although each basis function depends only on the single variable $\mathbf{x}_{i}$, it is formally regarded as a function defined over the entire input space [15]. During training, the MARS algorithm automatically determines the placement of knots and selects the hinge functions that contribute most to improving regression accuracy while ensuring generalization. The resulting regression model takes the form
@@ II-B Online Feedback Optimization
The OFO technique is based on the integration of the dynamics of a numerical optimization solver and the system dynamics to be controlled within a closed-loop structure. In this configuration, the solver is employed to update the control inputs for the system in response to the measurements obtained from the system. Note that this approach does not require an accurate plant model since is based on real-time measurements. Instead, it only requires the system’s input-output sensitivities, which can be obtained on site [26]. Therefore, the performance of the controller remains unaffected by model inaccuracies and is robust to disturbances and uncertainty, as these effects are implicit in the measurements.
The objective of OFO-based techniques is to steer the power system towards a steady-state optimal operating point. Therefore, under certain assumptions concerning the separation of time scales and the rapid mitigation of transients [24], the power system may be solely defined by its steady-state behavior. In this context, consider a set of nonlinear algebraic equations $\mathbf{f}(\cdot)$, as follows:
@@
The nonconvexity of the optimization problem (5a)-(5d) is addressed through the implementation of a projected gradient flow algorithm as presented in [13]. Starting from the control action $\mathbf{u}$ in the current iteration, the control algorithm proceeds to update the value ($\mathbf{u}^{+}$) using a projected gradient step method with a fixed step size [^1] A fixed step size algorithm must be used as an adverse consequence of not evaluating the set of equations $\mathbf{f}(\cdot)$ numerically. as follows:
@@
$\hat{\boldsymbol{\sigma}}(\mathbf{u,y}):=\arg\min_{\boldsymbol{\delta}\in\mathbb{R}^{p}}$ $\;\|\boldsymbol{\delta}+\gamma\mathbf{G}^{-1}\mathbf{F}^{\top}\nabla\phi(\mathbf{y,u})^{\top}\|_{\mathbf{G}}^{2}$ (8a)
@@ II-C Small-signal Stability Constrained Online Feedback Optimization
$\underset{\mathbf{u}}{\operatorname{minimize}}\quad$ $\gamma\cdot\phi(\mathbf{y},\mathbf{u})$
@@ III-A Test System
The simulation test bench is based on a modified version of the IEEE 9-bus system [1]. It represents a three-phase high-voltage transmission network consisting of nine buses, six transmission lines, three transformers, three load substations, and three generators. In contrast to the original system, which includes three SGs, two SGs are replaced with IBRs to emulate a scenario with high penetration of power electronics devices. Furthermore, to account for the impact of control strategies on system stability, one IBR is operated under GFM control while the other is operated under GFL control.
Figure 1 shows the one-line diagram of the system under analysis. The electrical lines of the system are considered symmetrical and are modeled using a $\pi$-model. The character of the lines is predominantly inductive, and all parameters are obtained from [1]. Power substations are modeled as constant impedance balanced loads (R-L shunts load model). Table I summarizes the main characteristics of the substations demand, including the load participation factors (i.e., the percentage of the total system demand allocated to each load), the load power factors, and the overall range of system demand. Concerning the generators, it is assumed that both SG and IBRs are equipped with a primary power source that allows them to operate effectively at any feasible operating point. In addition, all generators have a communication system with a centralized control layer that calculates their setpoints. At the local level, the SG is incorporated with a governor and an AVR to regulate its steady-state outputs to the values dispatched by the proposed control layer. Therefore, the setpoints for SGs are defined as the injected active power and the generated voltage amplitude, measured at the bus to which they are connected. The IBR connected to bus 1 is configured to operate in a GFM mode, and therefore indicated as $\text{GFM}_{1}$. The device is equipped with an internal current control loop and an external voltage control loop. In addition, it contains an active power droop control stage, employed for synchronization, and reactive power droop. The IBR connected to bus 3 is configured to operate in a GFL mode, and therefore indicated as $\text{GFL}_{3}$. The GFL control scheme comprises a Phase-Locked Loop (PLL), an inner current control loop, an outer control loop for active power fed by a frequency droop control, and an outer reactive power control fed by a voltage droop. Table II summarizes the main generator parameters, including their rated power, rated voltage, and the assumed operating active power ranges.